. The residents of an agrarian country spend all of their income on cauliflower, cucumbers, and carrots. In 2017, they produced 10 heads of cauliflower valued at Rs.200, 5 kgs of cucumbers for Rs.75, and 50 carrots for Rs.50. In 2018, they produced 5 heads of cauliflower for Rs.125, 8 kgs of cucumbers valued at Rs.120, and 50 carrots valued at Rs.100.
Calculate the price of each vegetable in each year. Also calculate nominal GDP and real GDP taking 2018 as the base year and GDP-deflator index.
b. Assuming that a common person consumes 2 heads of cauliflower, one kg. of cucumbers and 10 carrots in each year, calculate CPI taking 2018 as the base year.
c. What is the inflation rate between 2017-18 measured from GDP-deflator index and CPI?
1
Expert's answer
2020-10-29T07:05:43-0400
"Solution"

In 2017, the residents spent a total of "Rs.200" for "10" heads of cauliflower, so Per cauliflower"=Rs.20, Rs.75\ for" "5 kgs" of cucumber, so per cucumber"=Rs.15" and "Rs.50\ for\ 50\ carrots, so\ per\ carrot=1"


In 2018, they spent "Rs.125" for "5" heads of cauliflower, so per cauliflower"=Rs.2.5, for\ Rs.120\ cucumbers\ for\ 8kgs, per\ cucumber=Rs.15\ and\ Rs.100\ for\ 50\ carrots,\ so\ per\ carrot=Rs.2"



Therefore using 2017 as the base year, the CPI for 2017 is "100" and the CPI for 2018 is "\frac{345}{325}=1.06154\times100\\
=106"

The inflation rate in 2017-2018 using the CPI "=(106-100)\times100\%\\
=6\%"